# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v4

@@ 1 Introduction
The $s$-concavity assumption of $\psi_{i}$ also arises naturally as a generalization of the commonly used log-concavity (recovered when $s=0$), and plays a crucial role in guaranteeing convergence analysis towards equilibrium (see Section 3.1). We emphasize that $s$-concavity, being a higher-order smoothness constraint than monotonicity, drives the convergence rate results.
@@ 2 Related Literature and Contributions
By Equation 1, we have $N$ single-index models $\mathbb{E}(y_{i}\mid\mathbf{p})=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle),\ i\in\mathcal{N}=\{1,2,\dots,N\},$ with unknown $\bm{\theta}_{i}\in\mathbb{R}^{N}$ and unknown links $\psi_{i}$. We assume $\psi_{i}$ is nondecreasing, which makes demand $\lambda_{i}$ nonincreasing in its own price $p_{i}$, a standard assumption in the economics literature (birge2024interfere; li2024lego). Monotonicity also permits fully data-driven, tuning-free nonparametric estimators (balabdaoui2019least). The monotone SIM is identifiable under mild conditions: $\|\bm{\theta}_{i}\|_{2}=1$, and $\mathbf{p}$ has a strictly positive density on its domain (see balabdaoui2019least). balabdaoui2019least also propose a joint estimator $(\widehat{\bm{\theta}}_{i,n},\widehat{\psi}_{i,n})$ with $L^{2}$ error of order $n^{-1/3}$ for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$, and study the normalized linear estimator
@@
We propose a sequential price competition online algorithm that provides sublinear regret. The first step to achieve this goal is to establish the existence and uniqueness of a Nash Equilibrium $\mathbf{p^{*}}$, which is the fixed point of an operator $\bm{\Gamma}$ (see Section 3.1 for more details). A necessary condition for its existence is that each seller’s virtual valuation
@@
Our key observation links the NE and shape constraints as follows: $\varphi_{i}^{\prime}(u)\geq c_{i}$ iff $\psi_{i}$ is $(c_{i}-1)$-concave (Proposition 3.5; see the section below for definition of $s$-concavity), with log-concavity recovered as a special case when $c_{i}=1$. This connection allows us to estimate $\psi_{i}$ via nonparametric least squares under $s$-concavity – yielding a fully data-driven, tuning-free estimation procedure.
@@
Different from previous results, we study the convergence rate for the more general class $\mathcal{F}_{h}$ (see Appendix H for details) that contains functions of the form $\psi=h\circ\phi:\mathcal{U}\to\mathbb{R}$, where $h:\mathbb{R}\rightarrow(0,\infty)$ is a known increasing function in $\mathrm{C}^{2}$ and $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is an unknown concave function. Specifically, we consider i.i.d. observations $(U_{i},Y_{i})\sim(U,Y)$ for $i=1,2,\dots,n$ where $Y_{1},\dots,Y_{n}$ are noisy representations of the mean function $\psi_{0}(u)=\mathbb{E}\left(Y|U=u\right)=h(\phi_{0}(u))$ and $U_{1}\leq U_{2}\leq\dots\leq U_{n}$ are contained in $\mathcal{U}$. We study the uniform convergence rate of $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$, where $\widehat{\phi}_{n}$ is the LSE
@@ 2.1 Summary of Key Contributions
$s$-concave mean demand functions. In many existing works, necessary conditions for the existence of the NE are derived by assuming that the virtual valuation function defined in Equation 3 is increasing, often by invoking log-concavity of the mean demand function. We show how all such assumptions can be cast under the more general framework of $s$-concavity. This reformulation allows the development of a fully data-driven, tuning-parameter-free algorithm using shape constraints.
Regret upper bound and convergence to equilibrium. We establish an upper bound on the total expected regret and analyze the convergence to the NE for a general exploration length of order $\tau\propto T^{\xi}$ for $\xi\in(0,1)$ (Theorem 5.4). Our results reveal the existence of an optimal choice of $\xi$ that minimizes the total expected regret for each seller, leading to a regret of order $\widetilde{\mathcal{O}}(N^{3/2}T^{5/7})$, where $\widetilde{\mathcal{O}}$ excludes log factors (Remark 5.6). Moreover, we show that by the end of the selling horizon, the joint prices set by sellers converge to Nash equilibrium prices at a rate of $\widetilde{\mathcal{O}}(N^{3/4}T^{-1/7})$.
Concentration inequality for $s$-concave regression functions in the supremum norm. Our work involves establishing a concentration inequality for the nonparametric LSE, under the supremum norm, for a large class of shape constraints that includes $s$-concavity (Appendix H). As a minor contribution, we derive a concentration inequality for the parametric component $\bm{\theta}_{i}$ of the monotone single index model (Proposition D.3), while previous results show convergence in probability or distribution (balabdaoui2019least).
@@ 3 Problem Formulation
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in birge2024interfere; li2024lego. The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in 3.1.
@@
The constants $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in (11)), however, firms can estimate them easily from historical sales data and operational capacity limits; moreover loose bounds suffice and do not affect the rates in Theorem 5.4. The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (chmielewski1981elliptically; balabdaoui2019least). Note that Equation 6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$. Regarding $\bm{\gamma}_{i}$, while many applications assume $\gamma_{ij}>0$, this restriction is not required for our algorithm or theoretical guarantees - our results remain valid when $\gamma_{ij}>0$. We allow $\gamma_{ij}$ to be negative to capture settings with negative cross-effects, such as vertically differentiated products, where raising a competitor’s price may reduce my demand if consumers perceive their product as higher quality.
@@
where $\bm{\Gamma}$ is called Best Response Operator, and its components $\Gamma_{i}$ are defined in (7).
@@ 3.1 Main Assumptions
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$, were $\mathbf{p}^{\star}$ is a NE (see (37) for a detailed bound). By the triangle inequality, we have
@@
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients: $(i)$ Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically; $(ii)$ Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$. Putting these together, we obtain (see (G.2) for a formal inequality):
@@
By twice differentiability of $\psi_{i}$, 3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by 3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
3.2 guarantees that the best-response map $\Gamma_{i}$ in (7) is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the existence of the Nash equilibrium, in line with standard results in competitive games (scutari2014real; li2024lego; tsekrekos2024variational). Instances where 3.2 is satisfied include linear demand models (li2024lego), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section I.1. The next result is an immediate consequence of Theorem 3, in scutari2014real.
@@
Under assumptions (3.1) and (3.2), there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in (8).
By 3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma 3.7 for the derivation)
@@
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce 3.4, which makes the $\varphi_{i}$’s invertible.
@@
3.4 covers the linear demand model by li2024lego for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in fan2024policy, Assumption 1 in chen2018robust, Equation 1 in cole2014sample and golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of 3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints (the proof is provided in Appendix B). Examples of $s$-concave functions can be found in Appendix L.
@@
3.6 generalizes Assumption 2 in li2024lego (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in kachani2007modeling. The proof of the following result is relegated to Appendix C. In practice, 3.6 states that the influence of competitors’ prices on seller $i$’s optimal response is sufficiently small relative to the sensitivity to its own price.
@@
In this case 3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.

### v5

@@ 1 Introduction
The $s$-concavity assumption of $\psi_{i}$ also arises naturally as a generalization of the commonly used log-concavity (recovered when $s=0$), and plays a crucial role in guaranteeing convergence analysis towards equilibrium (see Section˜3.1). We emphasize that $s$-concavity, being a higher-order smoothness constraint than monotonicity, drives the convergence rate results.
@@ 2 Related Literature and Contributions
By Equation˜1, we have $N$ single-index models $\mathbb{E}(y_{i}\mid\mathbf{p})=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}\rangle),\ i\in\mathcal{N}=\{1,2,\dots,N\},$ with unknown $\bm{\theta}_{i}\in\mathbb{R}^{N}$ and unknown links $\psi_{i}$. We assume $\psi_{i}$ is nondecreasing, which makes demand $\lambda_{i}$ nonincreasing in its own price $p_{i}$, a standard assumption in the economics literature (birge2024interfere; li2024lego). Monotonicity also permits fully data-driven, tuning-free nonparametric estimators (balabdaoui2019least). The monotone SIM is identifiable under mild conditions: $\|\bm{\theta}_{i}\|_{2}=1$, and $\mathbf{p}$ has a strictly positive density on its domain (see balabdaoui2019least). balabdaoui2019least also propose a joint estimator $(\widehat{\bm{\theta}}_{i,n},\widehat{\psi}_{i,n})$ with $L^{2}$ error of order $n^{-1/3}$ for $\widehat{\psi}_{i}(\langle\widehat{\bm{\theta}}_{i},\cdot\rangle)$, and study the normalized linear estimator
@@
We propose a sequential price competition online algorithm that provides sublinear regret. The first step to achieve this goal is to establish the existence and uniqueness of a Nash Equilibrium $\mathbf{p^{*}}$, which is the fixed point of an operator $\bm{\Gamma}$ (see Section˜3.1 for more details). A necessary condition for its existence is that each seller’s virtual valuation
@@
Our key observation links the NE and shape constraints as follows: $\varphi_{i}^{\prime}(u)\geq c_{i}$ iff $\psi_{i}$ is $(c_{i}-1)$-concave (Proposition˜3.5; see the section below for definition of $s$-concavity), with log-concavity recovered as a special case when $c_{i}=1$. This connection allows us to estimate $\psi_{i}$ via nonparametric least squares under $s$-concavity – yielding a fully data-driven, tuning-free estimation procedure.
@@
Different from previous results, we study the convergence rate for the more general class $\mathcal{F}_{h}$ (see Appendix˜H for details) that contains functions of the form $\psi=h\circ\phi:\mathcal{U}\to\mathbb{R}$, where $h:\mathbb{R}\rightarrow(0,\infty)$ is a known increasing function in $\mathrm{C}^{2}$ and $\phi:\mathcal{U}\rightarrow\mathbb{R}$ is an unknown concave function. Specifically, we consider i.i.d. observations $(U_{i},Y_{i})\sim(U,Y)$ for $i=1,2,\dots,n$ where $Y_{1},\dots,Y_{n}$ are noisy representations of the mean function $\psi_{0}(u)=\mathbb{E}\left(Y|U=u\right)=h(\phi_{0}(u))$ and $U_{1}\leq U_{2}\leq\dots\leq U_{n}$ are contained in $\mathcal{U}$. We study the uniform convergence rate of $\widehat{\psi}_{n}=h\circ\widehat{\phi}_{n}$, where $\widehat{\phi}_{n}$ is the LSE
@@ 2.1 Summary of Key Contributions
$s$-concave mean demand functions. In many existing works, necessary conditions for the existence of the NE are derived by assuming that the virtual valuation function defined in Equation˜3 is increasing, often by invoking log-concavity of the mean demand function. We show how all such assumptions can be cast under the more general framework of $s$-concavity. This reformulation allows the development of a fully data-driven, tuning-parameter-free algorithm using shape constraints.
Regret upper bound and convergence to equilibrium. We establish an upper bound on the total expected regret and analyze the convergence to the NE for a general exploration length of order $\tau\propto T^{\xi}$ for $\xi\in(0,1)$ (Theorem˜5.4). Our results reveal the existence of an optimal choice of $\xi$ that minimizes the total expected regret for each seller, leading to a regret of order $\widetilde{\mathcal{O}}(N^{3/2}T^{5/7})$, where $\widetilde{\mathcal{O}}$ excludes log factors (Remark˜5.6). Moreover, we show that by the end of the selling horizon, the joint prices set by sellers converge to Nash equilibrium prices at a rate of $\widetilde{\mathcal{O}}(N^{3/4}T^{-1/7})$.
Concentration inequality for $s$-concave regression functions in the supremum norm. Our work involves establishing a concentration inequality for the nonparametric LSE, under the supremum norm, for a large class of shape constraints that includes $s$-concavity (Appendix˜H). As a minor contribution, we derive a concentration inequality for the parametric component $\bm{\theta}_{i}$ of the monotone single index model (Proposition˜D.3), while previous results show convergence in probability or distribution (balabdaoui2019least).
@@ 3 Problem Formulation
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in birge2024interfere; li2024lego. The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in ˜3.1.
@@
The constants $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in equation 11), however, firms can estimate them easily from historical sales data and operational capacity limits; moreover loose bounds suffice and do not affect the rates in Theorem˜5.4. The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (chmielewski1981elliptically; balabdaoui2019least). Note that Equation˜6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$. Regarding $\bm{\gamma}_{i}$, while many applications assume $\gamma_{ij}>0$, this restriction is not required for our algorithm or theoretical guarantees - our results remain valid when $\gamma_{ij}>0$. We allow $\gamma_{ij}$ to be negative to capture settings with negative cross-effects, such as vertically differentiated products, where raising a competitor’s price may reduce my demand if consumers perceive their product as higher quality.
@@
where $\bm{\Gamma}$ is called Best Response Operator, and its components $\Gamma_{i}$ are defined in equation 7.
@@ 3.1 Main Assumptions
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$, were $\mathbf{p}^{\star}$ is a NE (see equation 37 for a detailed bound). By the triangle inequality, we have
@@
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients: $(i)$ Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically; $(ii)$ Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$. Putting these together, we obtain (see equation G.2 for a formal inequality):
@@
By twice differentiability of $\psi_{i}$, ˜3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by ˜3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
˜3.2 guarantees that the best-response map $\Gamma_{i}$ in equation 7 is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the existence of the Nash equilibrium, in line with standard results in competitive games (scutari2014real; li2024lego; tsekrekos2024variational). Instances where ˜3.2 is satisfied include linear demand models (li2024lego), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section˜I.1. The next result is an immediate consequence of Theorem 3, in scutari2014real.
@@
Under assumptions equation 3.1 and equation 3.2, there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in equation 8.
By ˜3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma˜3.7 for the derivation)
@@
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce ˜3.4, which makes the $\varphi_{i}$’s invertible.
@@
˜3.4 covers the linear demand model by li2024lego for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in fan2024policy, Assumption 1 in chen2018robust, Equation 1 in cole2014sample and golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of ˜3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints (the proof is provided in Appendix˜B). Examples of $s$-concave functions can be found in Appendix˜L.
@@
˜3.6 generalizes Assumption 2 in li2024lego (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in kachani2007modeling. The proof of the following result is relegated to Appendix˜C. In practice, ˜3.6 states that the influence of competitors’ prices on seller $i$’s optimal response is sufficiently small relative to the sensitivity to its own price.
@@
In this case ˜3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.